OpenStudy (anonymous):

How to prove, cot 4x (sin 5x + sin 3x) = cot x (sin 5x - sin 3x ) ?

OpenStudy (raden):

these identities might work sinA + sinB = 2sin1/2(A+B)cos1/2(A-B) sinA - sinB = 2cos1/2(A+B)sin1/2(A-B) and cot = cos/sin

OpenStudy (anonymous):

Ok Friend, I know these equations. But, how can i prove this?

OpenStudy (raden):

for LHS : cot 4x (sin 5x + sin 3x) = (cot4x/sin4x) (2sin1/2(5x+3x)cos1/2(5x-3x)) = (cot4x/sin4x) (2sin4xcosx) cancel the sin4x's = 2cos4xcosx now, do same idea for the RHS

OpenStudy (raden):

you have to show that LHS = RHS

OpenStudy (anonymous):

:) ok. thank you